OpenStudy (anonymous):

I need help with this (8+√11) ( 8-√11)

OpenStudy (anonymous):

This is a difference of squares

OpenStudy (anonymous):

\[x^2-y^2 = (x-y)(x+y)\]

jhonyy9 (jhonyy9):

64-11=53

OpenStudy (anonymous):

the answer is 53

OpenStudy (anonymous):

FOIL it 1st Outside Inside Last

OpenStudy (anonymous):

Wow this is a great website for flunkies to come have their homework done for them... :)